Sound exits a diffraction horn loudspeaker through a rectangular opening like a small doorway. Such a loudspeaker is mounted outside on a pole. In winter, when the temperature is 273 K, the diffraction angle θ has a value of 17o. What is the diffraction angle for the same sound on a summer day when the temperature is 311 K?
Answer #1

We know that diffraction angle is given by:

sin heta = lambda/D

lambda = wavelength of sound = v/f

D = width of the rectangular opening

v = speed of sound

f = frequency, So

sin heta = v/(f*D)

Now Assuming air is an ideal gas, speed of sound will be given by:

v = sqrt (gammakT/m)

gamma = Cp/Cv = ratio of specific heat capacities

k = Boltzmann constant,

T = temperature

m = average mass of the molecules of air

So,

sin heta = v/(f*D) = [1/(f*D)]*sqrt (gammakT/m)

sin heta = mf D2

Now Since gamma, k, m, f, and D are constant, So

sin heta propto sqrt T

hetas = diffraction angle in summer = ?

hetaw = diffraction angle in winter = 17 deg

Ts = Temperature in summer = 311 K

Tw = Temperature in winter = 273 K

So,

sine sína, Tr Au

{sin heta_{s}} = {sin heta_{w}}sqrt {rac{T_{s}}{T_{w}}}

sino,-sin 17-V273

sine 0.312

0, -arcsin(0.312)

e, 18.2

Diffraction angle in summer = e, 18.2

Please Upvote.

Comment below if you have any query.